The following question relates to an oligopoly market where the industry demand curve is P = 100 - Q. Derive the reaction curve for a Cournot duopolist where the industry demand curve is as stated above and the MC of production is zero.
Answer #1

consnot duopody: P=100-Q Q=aut Qir Muzzo Firm I TRI = PQI = [100-(Q+B)]Q, -1002, -2,2-2,22 MR, =dTRI 7Q1 = 100(1) ~QQ, at 22

The ={100 - 21+22] Q2 TR= 1000 - 2,2₂ – Q² MR = dTR₂ Laz MR2 = 100 (1) - Qi(1) -2Q2 MR2 = 100-Q, -209 profit-maximizing condi